Find the angles which is equal to its complement.
 Find the angles which is equal to its complement.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the measure of the required angle be x^o .
We know that, sum of measures of complementary angle pair is 90^o .
Then,
= x + x = 90^o
= 2x = 90^o
= x = 90/2
= x = 45^o
Hence, the required angle measures is 45^o .

Thanks